Find the roots of the quadratic equations using the quadratic formula. Express the roots, if they are not rational, in simplest radicl form.

4x^2 - 12x - 7 = 0

Answer by fractalier(2101)

4x^2 - 12x - 7 = 0 yields
x = (12 ± sqrt(144+112)) / 8
x = (12 ± sqrt(256)) / 8
x = (12 ± 16) / 8
x = 28 / 8 = 7/2 and
x = -4 / 8 = -1/2